OpenStudy (anonymous):

The human body can survive an acceleration of less than 250 m/s2. If a person is involved in a car accident with an initial speed of 120 km/hr and is protected by an airbag, over what distance must the airbag function in order for the person to survive the crash? _ What formulas am i going to use? Help me please. Or can someone give me an explanation? Thank u very much

OpenStudy (zephyr141):

use kinematics\[v^2=v_0^2+2ad\]solve for d. final velocity is going to be zero. initial velocity is given and d is given.

OpenStudy (zephyr141):

oh right. you might have to do some conversions from kilometers to meters

OpenStudy (zephyr141):

and hours to seconds.

OpenStudy (anonymous):

@zephyr141 yea so i got 33.33 m/s, plus i substituted it to the equation, Vo = 120 km/hr = 33.33 m/s Vf = 0 m/s a = 250 m/s^2 Vf^2 = Vo^2 + 2ad d = (Vf^2-Vo^2)/2a i calculated d, and i got -2.22 m

OpenStudy (zephyr141):

the person is slowing down. what makes a person slow down? a negative acceleration. so put in a negative sign there and you'll get the same thing but positive.

OpenStudy (anonymous):

@zephyr141 ah!! that!! yes okay, 2.22m :) got it, thanks!!

OpenStudy (zephyr141):

nice!